# High Resolution Millimeter Wave Imaging Based on FMCW Radar Systems at W-Band

### v2

@@ VII-A Image Formation
In this subsection, we present the result of applying the back-projection algorithm to the data-set collected from a bike in front of the imaging system.
@@
Figure 9: The result of applying the back-projection algorithm to the experimental data gathered from the bike.
Fig. 9 shows the result of applying the back-projection algorithm, given in (3), to the experimental data gathered from the bike.
@@
In Fig. 11, we have shown the point-cloud for the reconstructed image shown in Fig. 9. Fig. 11 includes the information about the intensity of the point reflectors
@@
In Fig. 12 we have shown all the point reflectors with the same intensity level.
@@ VII-B Apodization
In this sub-section, the result of the dual apodization and complex dual apodization methods are presented. We have used rectangular and Hamming window to perform the apodization method. The rectangular window represents the image without windowing. It is important to mention that multiple windows can be used for apodization which create multi-layer algorithm for side-lobe reduction while preserving the resolution limit of the system.
The dual apodization method is based on the absolute value of the reconstructed image and provides reasonable result. However, the complex dual apodization works on the real and imaginary part of the reconstructed image and yields an incredibly better result.
@@
Finally, Fig. 13-(c) depicts the result of applying the complex dual apodization technique. As it is clear from Fig. 13-(c), the result of the complex dual apodization is by far better than that of the dual apodization. Of course, the computational complexity of the complex dual apodization is slightly higher.
@@ VII-C Image De-noising
In this sub-section, we implement the modified Perona-Malik method, given in (LABEL:denoising_opt_EL), directly in Python to remove the noise from the reconstructed images.
The result of the de-noised image has been shown in Fig. 14. In order for the effect of the de-noising process to be seen clearly, we have shown all the images with 50 dB dynamic range.
@@
As can be seen from Fig. 14-(b) the de-noising algorithm provides a remarkable result. The fine structures of the target can be uncovered from the noise while the edges have been preserved. The result shown in Fig. 14-(b) has been obtained by applying (LABEL:denoising_opt_EL) to the windowed image shown in Fig. 14-(a). In Fig. 14-(a), we have used Hamming window in both the range and azimuth directions.
@@
VII-D Sub-Array-Based Processing
This subsection discusses the sub-array-based processing results. Fig. 15 shows remarkable results for implementing the sub-array-based processing.
We have partitioned the whole array into sub-arrays with length 10 and with 9 elements overlap between them. Subsequently, we use all these sub-arrays to perform the final stage of the beam-forming process and obtain the image.
In Fig. 15-(a), we have steered the beam toward left by 15 degree. In Fig. 15-(b), we are illuminating the bore-sight and in Fig. 15-(c), we have directed the beam toward the right by 15 degree.
By performing sub-array-based processing with sub-arrays of length equal to 10, we can achieve approximately 3 dB gain which is impressive. The reason that we can not obtain 10 dB gain, which is the real number for creating sub-arrays with the length equal to 10, is that when we create the sub-arrays, the overlap between them will force the noise of the system to become correlated. As a result, when we perform the final stage of the beam-forming process to generate the final image we will be deprived of the full array gain because of the correlation between the noise samples.
The results that we have presented in Fig. 15 are based on 50 dB dynamic range and in gray scale, thus, we can compare them with the reconstructed image shown in Fig. 14-(a).
The complete analysis of the gain of the sub-array-based processing and its dependency on the length of the sub-arrays, will deviate us from the main concept of this paper and will be the subject of our future work.
It should be noted that, the gain that we obtain by using the sub-array-based processing technique, which we have presented in this section, is exclusively the gain that is provided by the signal processing unit and that is the reason that the sub-array-based method is considered to be highly valuable.
However, as we mentioned before and as it is clear form the results presented in Fig. 15, when we increase the number of elements per each sub-array the overall field of view of the system becomes narrower. In other words, achieving higher gain using the sub-array-based approach comes at the cost of narrowing the field of view of the imagery system. As a result, to be able to perform imaging for the entire field of view of the imagery system, we need to steer the beam similar to what we have done in Fig. 15.
@@ VII-D Sub-Array-Based Processing
Figure 15: The result of sub-array-based processing. Sub-arrays of length equal to 10 have been created and all the sub-beams have been directed towards the location of the bike. Finally,all the elements of the array have been used to perform the beam-forming, a) steering the beam toward left by 15 deg, b) steering the beam toward the center c) steering the beam toward right by 15 deg.
VII-E Multi-Spectral-Based Imaging
The result of multi-spectral-based imaging based on 3 sub-bands has been shown in Fig. 16. The result is based on 3 separate images, one image per each sub-band. Next, they are assigned to red, green, and blue colors and the final image is created. For better visualization, we have presented $1-|\text{img}|$ which implies that if the target has the same response over the entire bandwidth, it will be shown in black color.
In order to increase the visibility of the image, we have performed contrast enhancement as well [26]. We have applied the de-noising algorithm given in (LABEL:denoising_opt_EL) to the image prior to partitioning the data into 3 sub-bands.
In Fig. 16, the target with darker color shows the same response over the entire bandwidth of the system. In other words, its reflective coefficient has less dependency on the range of frequencies that the system is operating over.
Refer to captionRange [m]Cross Range [m]
The multi-spectral-based imaging can be used for material characterization and this will be the subject for our future work.
@@ VII-F More Results
In this subsection, we present more experimental results from targets such as bike as well as car at near and far distances from the imagery system. It should be noted that, for all the experimental results presented in this subsection, we have used the back-projection algorithm, given in (3), to reconstruct the images. Also, we have used the de-noising algorithm given in (LABEL:denoising_opt_EL) to de-noise the images.
We have performed a test in which a car and two bikes have been located in front of the radar system. Fig. 17-(a) illustrates the experimental set-up. In Fig. 17-(b), we have presented the reconstructed image. From Fig. 17-(b), we can see all the details of all targets present in the radar’s field of view.
@@
We have conducted a different experiment which includes 2 cars in front of the imagery system. Fig. 18-(a) illustrates the experimental set-up. As can be seen from Fig. 18-(a) , the data has been collected during night time. In Fig. 18-(b), we have shown the reconstructed image. From Fig. 18-(b), we can see the details of both cars such as the bumper, headlights, license plate, and side mirror.
We have conducted an additional test. The data for this scenario has been collected in rainy weather. Fig. 19-(a) illustrates the experimental set-up for the bike situated on the ground in front of the car. As we mentioned, the data has been collected in a rainy morning in order to test the ability of the imagery system to create a high resolution image from the targets even in rainy condition. As can be seen from Fig. 19-(a), the ground, the bike, and the car are all wet and rain is coming during the period that the data is being collected. Fig. 19-(b) shows the reconstructed image.
@@
To further demonstrate the all-weather capability of the imagery system, we have conducted a test in a foggy environment. The data for this scenario has been collected in dense fog. Fig. 20-(a) shows the experimental set-up for the bike situated on the ground in front of the imagery system in clear weather condition and Fig. 20-(b) presents the reconstructed image from the scene. The set-up is inside the garage and we have then created a thick fogy situation. In Fig. 21-(a), the test set-up in the fog has been shown. The thick layer of fog makes it impossible for the bike to be seen in the optical image presented by Fig. 21-(a). In Fig. 21-(b), we have presented the reconstructed image form the fogy test set-up. It is clear form In Fig. 21-(b) that even in thick fogy situation the imagery system is capable of producing high resolution image form the target. In fact the reconstructed images under no-fog and foggy conditions, presented in Fig. 20-(b) and Fig. 21-(b), respectively, are identical.
@@
Figure 21: a) the experimental set-up presented in Fig. 20-(a), the data has been collected in dense fog, b) the reconstructed image.
@@
It should be noted that, for this experiment we have used horn antennas on both the transmit and receive sides in order to increase the maximum range of the imagery system. Consequently, the field of view of the imagery system has been reduced and in the reconstructed image, which has been shown in Fig. 23-(b), we can only see the bike.
@@ VIII Conclusion
In this paper, we presented our custom-made FMCW radar system which operates at millimeter wave frequencies. We described the functionalities of the system in detail. We further, presented the imaging algorithm and several signal processing techniques for image formation and image post processing. Finally, we presented all the results based on the experimental data gathered from the imagery system. The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost millimeter wave systems such as the one presented in this work.

### v3

@@ VII-A Image Formation
In this subsection, we present the result of applying the back-projection algorithm to the data-set collected from a bike in front of the imagery system.
@@
Figure 9: The result of applying the back-projection algorithm, based on (3), to the experimental data gathered from the bike.
Fig. 9 shows the result of applying the back-projection algorithm, based on (3), to the experimental data gathered from the bike.
@@
In Fig. 11, we have shown the point-cloud for the reconstructed image shown in Fig. 9.
Fig. 11 includes the information about the intensity of the point reflectors.
@@
In Fig. 12, we have shown all the point reflectors with the same intensity level.
@@ VII-B Apodization
In this sub-section, the result of the dual apodization and complex dual apodization methods are presented. We have used rectangular and Hamming window to perform the apodization method. The rectangular window represents the image without windowing. It is important to mention that multiple windows can be used for apodization which create multi-layer algorithm for side-lobe reduction while the resolution limit of the system is preserved.
The dual apodization method is based on the absolute value of the reconstructed image and provides a reasonable result. However, the complex dual apodization operates on the real and imaginary parts of the reconstructed image separately and yields an incredibly better result with significantly lower side-lob levels.
@@
Finally, Fig. 13-(c) depicts the result of applying the complex dual apodization technique. As it is clear from Fig. 13-(c), the result of the complex dual apodization is by far better than that of the dual apodization. It should also be noted that, the computational complexity of the complex dual apodization is slightly higher than that of the dual apodization.
@@ VII-C Image De-noising
In this sub-section, we implement the modified Perona-Malik method, given in (LABEL:denoising_opt_EL), directly in Python to remove the noise from the reconstructed images. The result of the de-noised image has been shown in Fig. 14. In order for the effect of the de-noising process to be seen clearly, we have shown all the images with 50 dB dynamic range.
@@
As can be seen from Fig. 14-(b), the de-noising algorithm provides a remarkable result. The fine structures of the target can be uncovered from the noise while the edges have been preserved. The result shown in Fig. 14-(b), has been obtained by applying (LABEL:denoising_opt_EL) to the windowed image shown in Fig. 14-(a). In Fig. 14-(a), we have used Hamming window in both the range and azimuth directions.
@@
VII-D Multi-Spectral-Based Imaging
The result of multi-spectral-based imaging based on 3 sub-bands has been shown in Fig. 15. The result is based on 3 separate images, one image per each sub-band. Then, they are assigned red, green, and blue colors and the final image is created. In order to increase the visibility of the image, we have performed contrast enhancement as well [25]. We have applied the de-noising algorithm given in (LABEL:denoising_opt_EL) to the image prior to partitioning the data into 3 sub-bands.
For better visualization, we have plotted $\rm 1-|img|$, in which $\rm|img|$ represents the absolute value of the normalized image. Therefore, in Fig. 15, a target with darker color shows the same response over the entire bandwidth of the system. In other words, its radar cross section has less dependency on the range of frequencies that the system is operating over.
Refer to captionRange [m]Cross Range [m]
As we mentioned before, the multi-spectral-based imaging can be used for material characterization which will be the subject of our future work.
VII-E Sub-array Processing
This subsection discusses the sub-array-based processing results. Fig. 16 shows remarkable results for implementing the sub-array-based processing with $m=10$, where $m$ is the length of the sub-arrays. We have partitioned the entire array into sub-arrays with length 10 and with 9 elements overlap between them. Subsequently, we use all these sub-arrays to perform the final stage of the beam-forming process and obtain the image. In Fig. 16-(a), we have steered the beam toward left by $\rm 15^{o}$. In Fig. 16-(b), we are illuminating the bore-sight and in Fig. 16-(c), we have directed the beam toward right by $\rm 15^{o}$.
By performing sub-array-based processing with sub-arrays of length equal to 10, we can achieve approximately 3 dB gain which is impressive. The reason that we can not obtain 10 dB gain, which is the real number for creating sub-arrays with the length equal to 10, is that when we create the sub-arrays, the overlap between them will force the noise of the system to become correlated. As a result, when we perform the final stage of the beam-forming process to generate the final image, we will be deprived of the full array gain because of the correlation between the noise samples. The results that we have presented in Fig. 16, are based on 50 dB dynamic range and have been presented in gray scale, thus, we can compare them with the reconstructed image shown in Fig. 14-(a). The complete analysis of the gain of the sub-array-based processing and its dependency on the length of the sub-arrays, will deviate us from the main subject of this paper and will be the subject of our future work. It should be noted that, the gain that we obtain by using the sub-array-based processing technique, which we have presented in this section, is exclusively the gain that is provided by the signal processing unit and that is the reason that the sub-array-based method is considered to be highly valuable.
However, as we mentioned before and as it is clear from the results presented in Fig. 16, when we increase the number of elements per each sub-array, the overall field of view of the system becomes narrower. In other words, achieving higher gain using the sub-array-based approach comes at the cost of narrowing the field of view of the imagery system. As a result, in order to be able to perform imaging for the entire field of view of the imagery system, we are required to steer the beam similar to what we have presented in Fig. 16.
@@ VII-E Sub-array Processing
Figure 16: The result of sub-array-based processing. Sub-arrays of length equal to 10 have been created and all the sub-beams have been directed towards the location of the bike. Finally, all the elements of the array have been used to perform the beam-forming, a) steering the beam toward left by $\rm 15^{o}$, b) steering the beam toward the center c) steering the beam toward right by $\rm 15^{o}$.
@@ VII-F More Results
In this subsection, we present more experimental results from targets such as bike as well as car at near and far distances from the imagery system. It should be noted that, for all the experimental results presented in this subsection, we have used the back-projection algorithm, based on (3), to reconstruct the images. Also, we have used the de-noising algorithm based on (LABEL:denoising_opt_EL) to de-noise the images.
We have performed a test in which a car and two bikes have been located in front of the radar system. Fig. 17-(a) illustrates the experimental set-up. In Fig. 17-(b), we have presented the reconstructed image. From Fig. 17-(b), we can see all the details of all targets that are present in the radar’s field of view.
@@
We have conducted a different experiment which includes 2 cars in front of the imagery system. Fig. 18-(a) illustrates the experimental set-up. As can be seen from Fig. 18-(a) , the data has been collected during night time. In Fig. 18-(b), we have shown the reconstructed image. From Fig. 18-(b), we can see the details of both cars such as the bumper, headlights, license plates, and side mirrors.
We have conducted an additional test. The data for this scenario has been collected in rainy weather. Fig. 19-(a) illustrates the experimental set-up for the bike situated on the ground in front of the car. As we mentioned, the data has been collected in a rainy morning in order to test the ability of the imagery system to create high resolution images from the targets even in rainy conditions. As can be seen from Fig. 19-(a), the ground, the bike, and the car are all wet and rain is coming during the period that the data is being collected. Fig. 19-(b) shows the reconstructed image.
@@
To further demonstrate the all-weather capability of the imagery system, we have conducted a test in a foggy environment. The data for this scenario has been collected in dense fog. Fig. 20-(a) shows the experimental set-up for the bike situated on the ground in front of the imagery system in clear weather condition and Fig. 20-(b) presents the reconstructed image from the scene. The set-up is inside the garage and we have then created a thick fogy situation. In Fig. 21-(a), the test set-up in the fog has been shown. The thick layer of fog makes it impossible for the bike to be seen in the optical image which has been presented in Fig. 21-(a). In Fig. 21-(b), we have illustrated the reconstructed image from the fogy test set-up. It is clear from Fig. 21-(b) that even in thick fogy situation the imagery system is capable of producing high resolution images from the targets. In fact, the reconstructed images under no-fog and foggy conditions, presented in Fig. 20-(b) and Fig. 21-(b), respectively, are identical.
@@
Figure 21: a) the experimental set-up presented in Fig. 20-(a), the data has been collected in a foggy condition, b) the reconstructed image.
@@
It is worth mentioning that, for this experiment we have used horn antennas on both the transmit and receive sides in order to increase the maximum range of the imagery system. Consequently, the field of view of the imagery system has been reduced and in the reconstructed image, which has been shown in Fig. 23-(b), we can only see the bike.
@@ VIII Conclusion
In this paper, we presented our wide-band custom-made FMCW radar system which operates at millimeter wave frequencies. We described the functionalities of the system in detail. We further, presented the imaging algorithm and several signal processing techniques for image formation and image post processing. At the end, we presented all the results based on the experimental data gathered from the imagery system.
The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost millimeter wave systems such as the one presented in this work.